# Perturbation-based Analysis of Compositional Data

### v1

@@ 3.2 Efficient estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. We postpone the theoretical analysis and a detailed description of cross-fitting to Section 5 and only provide a high-level overview here. All the presented estimators involve regressions that can be performed by arbitrary (sufficiently powerful) regression methods.
@@ 3.2.1 Nonparametric estimation of the binary perturbation effect
Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathbbm{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
@@
The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities). The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of average predictive effects (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature. The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ is related to the augmented inverse propensity weighted (AIPW) estimator and is given by
@@ 3.2.2 Nonparametric estimation of the directional perturbation effect
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023). Here, we use an alternative smoothing method which uses random forests and local polynomials. A full description of our smoothing method is provided in Section S3.3 of the supplementary material. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm 1). Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplement.
@@ 3.2.3 Partially linear estimation of perturbation effects
While the estimators in the previous sections make only mild assumption on the data generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions. One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
@@ 4 Numerical experiments
We now present three numerical experiments illustrating the usage and performance of our methods. In Section 4.1 we investigate the performance of the semiparametric estimators discussed in Section 3. In Section 4.2 we analyze the effect of diversity on income based on grouped data from the 1994 US census. In Section 4.3 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.3 where we used the c-lasso package (Simpson et al., 2021).
@@ 4.1 Semiparametric estimators
  • plugin: The naive plugin versions of the NPM estimates. In the binary effect case, the estimator in (15) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (16) without the term involving $\widehat{\rho}$.
@@
The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plugin is estimated using an estimator akin to the one given in (14). Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{g}$ and $\widehat{m}$ and NPM and NPM_no_x additionally on the score estimate $\widehat{\rho}$. For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction. For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplement, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
We consider four data generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$. In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathbbm{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$. We then consider the following four generative models
@@
    $L=U_{0}B(W_{1})+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
@@
Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedures over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$. PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting. NPM works in both settings but can be less robust for small sample sizes. Further, it can be seen that the plugin estimator and the estimators without cross-fitting do not lead to correct coverage in some of these settings.
@@
In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated. We see that the plugin estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$. In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the tests, e.g. for $n=1000$ and binary $L$. In the cases where cross-fitting is not necessary, it does not seem to negatively impact the performance of the test.
4.2 Compositional confounding in relationship between income and race
@@
In this section we provide an example based on real data that illustrates how failing to adjust for compositional confounding can lead to incorrect conclusions about effects of compositional variables. We consider the ‘Adult’ dataset available from the UCI machine learning repository (Becker and Kohavi, 1996). This dataset is based on an extract from the 1994 US Census database and consists of data on 48,842 individuals. The response variable in the data, compensation, is a binary variable where 00 and $1$ encode whether an individual makes below or above 50,000 USD per year, respectively. Additionally, we consider the demographic features age (numeric, between 17 and 90), education (categorical, with 16 categories), sex (binary, 00 encoding male and $1$ female) and race (categorical, with categories ’White’, ’Black’ and ’Other’).
@@
We now use this individual-level data to construct community-level datasets by artificially grouping individuals together. This allows us to construct settings, where there is a known dependence between the racial diversity of the communities and the response variable. As we show below, this illustrates the potential pitfalls of ignoring compositional structure in the data and how the approach based on perturbations solves these issues. More specifically, we group all individual-level observations into groups of around $50$ individuals. We then create a new response variable $Y\in[0,1]$ by averaging the individual-level compensation in each group. Similarly, we create a predictor variable $X\in\mathbb{R}\times[0,1]\times\{1,\dots,16\}$ by averaging age and sex and selecting the majority class of the categorical variable education (breaking ties randomly). Finally, we create a compositional predictor $Z\in\Delta^{2}$ by computing the proportions of ‘White’, ‘Black’ and ‘Other’ individuals in each group. To ensure, a (partially) known effect of increased racial diversity on compensation, we construct three categories of groups (0, 1 and 2). We construct these categories in two different ways:
@@
Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$). Depending on the color the dependence between $Y$ and the Gini coefficient changes.
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use minus the Gini coefficient $L=-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 5, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3. The naive estimator that conditions either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded (e.g., educational levels may well be correlated with race and average compensation in actual communities).
These pitfalls can be avoided with a perturbation-based analysis. In this setting, we propose to apply $\mathrm{CDI}_{\mathrm{Gini}}$ with and without conditioning on the additional covariates $X$. The results are also shown in Table 3. As desired one gets significant positive effects with this method as it is adapted to the simplex. Interestingly, for $\mathrm{CDI}_{\mathrm{Gini}}$ that adjusts for $X$ the effect disappears for the grouping based on education. This is expected because the association between diversity and compensation was constructed based on education which is also part of $X$.
@@
naive_diversity ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\pgfsys@color@rgb@stroke{0.83984375}{0.15234375}{0.15625}\pgfsys@color@rgb@fill{0.83984375}{0.15234375}{0.15625}\mathbf{-0.082}}$ $(-0.144,-0.019)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\pgfsys@color@rgb@stroke{0.83984375}{0.15234375}{0.15625}\pgfsys@color@rgb@fill{0.83984375}{0.15234375}{0.15625}\mathbf{-0.120}}$ $(-0.194,-0.046)$
@@
$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.233}}$ $(0.060,0.406)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.614}}$ $(0.429,0.799)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\pgfsys@color@rgb@stroke{0.171875}{0.62890625}{0.171875}\pgfsys@color@rgb@fill{0.171875}{0.62890625}{0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
Table 3: Table of estimates in the ‘Adult’ data experiment. Bold numbers indicate significance at a $5\%$ level, red indicates a negative effect, while green indicates a positive effect of diversity. The naive_diversity method indicates that either there is a significant negative or a highly uncertain effect of diversity on compensation despite the fact that the data is constructed with a positive effect of diversity. The CDI correctly identifies this effect.

### v2

@@ 3.2 Efficient estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. We postpone the theoretical analysis and a detailed description of cross-fitting to Section 5 and only provide a high-level overview here. All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
@@ 3.2.1 Nonparametric estimation of the binary perturbation effect
Consider a binary perturbation $\psi$ with corresponding perturbation effect $\lambda_{\psi}$ and recall from the arguments of Section 2 that with $L:=\mathds{1}_{\{Z=\psi(Z,1)\}}$ and $W:=\psi(Z,1)$, we have that
@@
The denominator and $\mathbb{E}[Y]$ are both easy to estimate efficiently using their empirical counterparts (the one-step correction term is 00 for these quantities). The term $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ in the numerator is trickier and is related to the notion of the average predictive effect (e.g., Robins et al., 1994; Chernozhukov et al., 2018; Kennedy, 2023) that has been analyzed extensively in the causal inference literature. The one-step corrected estimator for $\mathbb{E}\big{[}\mathbb{E}\left[Y\,|\,L=1,W\right]]$ corresponds to the augmented inverse propensity weighted (AIPW) estimator and is given by
@@ 3.2.2 Nonparametric estimation of the directional perturbation effect
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023). Here, we use an alternative smoothing method which uses random forests and local polynomials. A full description of our smoothing method is provided in Section S3.3 of the supplementary material. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm 1). Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplementary material.
@@ 3.2.3 Partially linear estimation of perturbation effects
While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions. One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
@@ 4 Numerical experiments
We now present three numerical experiments illustrating the usage and performance of our methods. In Section 4.1 we investigate the performance of the semiparametric estimators discussed in Section 3. In Section 4.2.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census. In Section 4.2.2, we analyze the effect of racial diversity on grades in New York schools. In Section 4.3 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.3 where we used the c-lasso package (Simpson et al., 2021).
@@ 4.1 Semiparametric estimators
  • plugin: The naive plug-in versions of the NPM estimates. In the binary effect case, the estimator in (15) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (16) without the term involving $\widehat{\rho}$.
@@
The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (14). Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$. For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction. For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplementary material, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
We consider four data-generating models each with varying sample size $n\in\{250,1000,4000\}$ and predictor dimension $d\in\{3,15,75\}$. In all settings, we let $W$ be uniformly distributed on $\Delta^{d-2}$, denote by $\mathrm{med}_{W_{1}}$ the population median of $W_{1}$ and $\sigma_{W_{1}}^{2}$ the variance of $W_{1}$ and define $B(W_{1}):=\mathds{1}_{\{W_{1}>\mathrm{med}_{W_{1}}\}}$. We then consider the following four generative models
@@
    $L=U_{0}(1-B(W_{1}))+U_{1}B(W_{1})\quad\text{and}\quad Y=\frac{7}{5}LB(W_{1})+\sigma_{W_{1}}^{-1}W_{1}+\varepsilon$
@@
Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$. PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting. NPM works in both settings but can be less robust for small sample sizes. Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.
@@
In all settings we see that the partially linear model works well when it is well-specified but can fail when the partially linear assumption is violated. We see that the plug-in estimator generally performs worse than the one-step estimator but this disparity is reduced for large $n$. In many cases cross-fitting is not a necessity, however, in some settings cross-fitting does positively impact the performance of the confidence intervals, e.g. for $n=1000$ and binary $L$. In the cases where cross-fitting is not necessary, it does not seem to negatively impact the coverage of the intervals.
4.2 Estimating effects of diversity on socio-economic outcomes
4.2.1 Effects of diversity on income in semi-synthetic US census data
@@
In this section we provide an example based on semi-synthetic data that illustrates how failing to adjust for compositional confounding can lead to incorrect conclusions about effects of compositional variables. We consider the ‘Adult’ dataset available from the UCI machine learning repository (Becker and Kohavi, 1996). This dataset is based on an extract from the 1994 US Census database and consists of data on 48,842 individuals. The response variable in the data, compensation, is a binary variable where 00 and $1$ encode whether an individual makes below or above 50,000 USD per year, respectively. Additionally, we consider the demographic features age (numeric, between 17 and 90), education (categorical, with 16 categories), sex (binary, 00 encoding male and $1$ female) and race (categorical, with categories ’White’, ’Black’ and ’Other’).
We now use this raw individual-level data to construct community-level datasets by artificially grouping individuals together. The grouping is deliberately constructed in a way such that diversity affects income and this relationship is confounded by the remaining variation of the racial composition. This semi-synthetic example is only intended as an illustration. A genuine analysis of this data would require knowledge of the true community structure, which is not available in the data extract we are using here. We emphasize that even if the individual-level data is available, one may still be interested in effects of the racial composition of the community an individual lives in on that individual’s income. For such an analysis, one can add the racial composition of an individual’s community as a variable and then perform the same analysis as here but now on the individual-level data while adjusting for individual-level confounding factors. Importantly, even for such an individual-level analysis, the simplex confounding can bias the results in the same way as we illustrate here. We provide an example of this bias occurring on the individual-level effect in Section S4.4 of the supplementary material.
@@
We group all individual-level observations into groups of around $50$ individuals. We then create a new response variable $Y\in[0,1]$ by averaging the individual-level compensation in each group. Similarly, we create a predictor variable $X\in\mathbb{R}\times[0,1]\times\{1,\dots,16\}$ by averaging age and sex and selecting the majority class of the categorical variable education (breaking ties randomly). Finally, we create a compositional predictor $Z\in\Delta^{2}$ by computing the proportions of ‘White’, ‘Black’ and ‘Other’ individuals in each group. To ensure, a (partially) known effect of increased racial diversity on compensation, we construct three categories of groups (0, 1 and 2). We construct these categories in two different ways:
@@
Bottom: Plot of average compensation per group ($Y$) against Gini coefficient colored by the value of $W=(z_{\mathrm{cen}}-Z)/\|z_{\mathrm{cen}}-Z\|_{1}$ ($W$ takes values in a $1$-dimensional subset of $\mathbb{R}^{3}$ and can be represented by a point on unit circle). Depending on the color the dependence between $Y$ and the Gini coefficient changes.
Based on the two constructed community-level datasets, we now compare a naive approach that neglects the simplex structure and is applied in the literature (e.g. Richard (2000); Richard et al. (2007); Lu et al. (2015); Paarlberg et al. (2018)) with the perturbation-based approach. The naive approach summarizes racial diversity in a single measure, here we use minus the Gini coefficient $L=-\frac{1}{2d}\sum_{j=1}^{d}\sum_{k=1}^{d}|Z^{j}-Z^{k}|$ as in Example 5, for which large values correspond to high diversity, and estimate associations between this measure and the average compensation $Y$. We call the method that regresses $Y$ on $L$ linearly naive_diversity. This estimator can easily be adapted to also adjust for the additional covariates $X$, the racial composition $Z$ or both under a partially linear model assumption. The point estimates together with asymptotic confidence intervals (based on an efficient partialling out estimator like the estimator in (3)) are provided in Table 3 where random forests are used for nuisance function estimation when including additional covariates. The naive estimators that condition either on the empty set or on $X$ both result in significant negative effects which by construction of the datasets is incorrect. Furthermore, additionally conditioning on $Z$, leads to an ill-conditioned estimator as $Z$ is now a perfect predictor of the diversity measure. This is also visible in the large resulting confidence intervals. While we here constructed the groups in a way to highlight how compositional confounding can be misleading, we believe that this type of confounding is likely also present in real-world settings as groupings are often confounded (e.g., educational levels may well be correlated with race and average compensation in actual communities).
These pitfalls can be avoided with a perturbation-based analysis. In this setting, we propose to apply $\mathrm{CDI}_{\mathrm{Gini}}$ with and without conditioning on the additional covariates $X$. The results are also shown in Table 3 again using random forests for nuisance function estimation. As desired one gets significant positive effects with this method as it is adapted to the simplex. Interestingly, for $\mathrm{CDI}_{\mathrm{Gini}}$ that adjusts for $X$ the effect disappears for the grouping based on education. This is expected because the association between diversity and compensation was constructed based on education which is also part of $X$.
@@
naive_diversity ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.082}}$ $(-0.144,-0.019)$ ${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\mathbf{-0.120}}$ $(-0.194,-0.046)$
@@
$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.233}}$ $(0.060,0.406)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.614}}$ $(0.429,0.799)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ $0.071$ $(-0.049,0.191)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.611}}$ $(0.455,0.768)$
Table 3: Table of estimates in the ‘Adult’ data experiment. Bold red numbers indicate a significant negative effect of diversity at a $5\%$ level, while bold green indicates a significant positive effect. The naive_diversity method indicates that either there is a significant negative or a highly uncertain effect of diversity on compensation despite the fact that the data is constructed with a positive effect of diversity. The CDI correctly identifies this effect.
4.2.2 Effects of diversity on grades in New York schools
In this section we analyze data from a 2018 Kaggle competition containing data on schools in New York (PASSNYC, 2018). The data (collected in 2016) contains average grades in English Language Arts (ELA) and in math for 1217 schools along with additional background information on the school such as a summary of the economic need of the students, percentage of students with English as second language, racial composition of the school and various indicators of school performance (for a full list of control variables and our pre-processing of these, see Section S4.5 of the supplementary material). We let $X$ denote all background variables except racial composition which we denote by $Z$. Race is given as whole percentages of ‘Asian’, ‘Black’, ‘Hispanic’ and ‘White’ students which we rescale so that $Z\in\Delta^{3}$. We are interested in the effect of diversity, again measured by the Gini coefficient, on the average grade of the schools (either ELA or math). We repeat the analysis in Section 4.2.1 on this real data and the results can be seen in Table 4.
The naive_diversity method indicates that there is a significant effect of diversity on grades even when controlling for additional covariates $X$. In contrast, the CDI estimates smaller positive effects for ELA and effects that cannot be distinguished from 00 for math. This discrepancy between the two methods indicates that there is a significant amount of compositional confounding in this data, which invalidates the marginal analysis.
Method Effect on ELA grades Effect on math grades
Estimate $95\%$ CI Estimate $95\%$ CI
naive_diversity ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{1.079}}$ $(0.934,1.223)$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{1.231}}$ $(1.040,1.422)$
[…]
$\mathrm{CDI}_{\mathrm{Gini}}$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.345}}$ $(0.105,0.585)$ $0.098$ $(-0.175,0.371)$
$\mathrm{CDI}_{\mathrm{Gini}}\,|\,X$ ${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\mathbf{0.163}}$ $(0.002,0.323)$ $0.047$ $(-0.167,0.261)$
Table 4: Table of estimates in the New York schools data experiment. Bold green numbers indicate a significant positive effect of diversity at a $5\%$ level. The naive_diversity method indicates that there is a significant effect of diversity on grades even when controlling for additional covariates $X$. The CDI (with or without controlling for $X$) generally estimates smaller positive effects or even effects that cannot be distinguished from 00, indicating a significant amount of compositional confounding in this data.